Car A leaves home at 40 m/h.Four hours later car B leaves from a house that is 700 miles away headed towards car A at 50 m/h.How long will it take before the two cars meet?
Answer by TimothyLamb(4379)

s = d / t
d = s * t
t = d / s
---
when b leaves a has already traveled:
d = 40 * 4 = 160 miles
so when b leaves, the distance between a and b is:
700 - 160 = 540 miles
---
the closing speed of the two cars is:
40 + 50 = 90 mph
---
when the cars meet they will have covered 540 miles between both cars, at a closing speed of 90 mph, so:
t = 540 / 90 = 6 hours
---
Answer:
the cars will meet:
6 hours after b leaves
10 hours after a leaves (because a left 4 hours earlier than b)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php